A compound containing carbon, hydrogen and oxygen has been analyzed. It contains 69.72% of carbon and 18.58% of oxygen. If the molecular weight of the compound is 86.13, what is its molecular formula?

NOTE: Write the compound in the form (CXHYOZ), where X, Y and Z are whole numbers.
Answer #1

100gm of compound contains

C = 69.72gm*1/12.01*mol/gm= 5.8

O = 18.58gm*1/16mol/gm = 1.16

H = 100-(69.72+18.58) = 11.7gm*1/1.01 mol/gm = 11.6

Mole ratio:

Dividing by smallest mol 1.16

C =5.8/1.16=5

O=1.16/1.16=1

H=11.6/1.16=10

So mol ratios give empirical formula C5OH10

Empirical formula weight = 5*12.01+1*16+10*1.01= 86.16 which is nearly equal to the molecular weight 86.13

So molecular formu

la is C5OH10